How would I find the hypotenuse of a triangle if a=5 and b=12?
It tells me the answer is 13 but I don't understand how..

Answer by Algebra2you(12)

Remember, the Pythagorean theorem states that a²+b²=c².
a and b are the legs. The order doesn't matter. c is the hypotenuse, or the side opposite the right angle. It is also the longest side in a triangle.
So let's substitute a=5 and b=12 into the Pythagorean theorem and solve for c, the hypotenuse.
5²+12²=c²
25+144=c² Simplify squares.
169=c² Add like terms.
√169=√c² Take the square root of both sides to isolate the c.
13=c Simplify
So c, or the hypotenuse, is 13 :)